5. Evaluate where D is the upper solid hemisphere 2y2+ z2 < 4, z 2 0.
Answer #1

Evalute ý ar where D is the upper solid hemisphere ants 34,820. √ 4-z-zz J y dx dy dz 8=0 y = -√4-82 x=-14-y - 22 13 (0,0,1)3) sino caso do dz det y = √ 4-zr simo. t.dy = √4-87 cost do. dz / 4 sino Cose do -- 6 | 8 | 6 | 10 sinzo do J(-2) 448 44-2 5